# FSOS-AMC: Few-Shot Open-Set Learning for Automatic Modulation Classification

### v1

@@ Abstract
Automatic modulation classification (AMC) is essential for the advancement and efficiency of future wireless communication networks. Deep learning (DL)-based AMC frameworks have garnered extensive attention for their impressive classification performance. However, existing DL-based AMC frameworks rely on two assumptions, large-scale training data and the same class pool between the training and testing data, which are not suitable for few-shot and open-set scenarios. To address this issue, a novel few-shot open-set automatic modulation classification (FSOS-AMC) framework is proposed by exploiting a multi-scale attention network, meta-prototype training, and a modular open-set classifier. The multi-scale attention network is used to extract the features from the input signal, the meta-prototype training is adopted to train the feature extractor and the modular open-set classifier can be utilized to classify the testing data into one of the known modulations or potential unknown modulations. Extensive simulation results demonstrate that the proposed FSOS-AMC framework can achieve higher classification accuracy than the state-of-the-art methods for known modulations and unknown modulations in terms of accuracy and area under the receiver operating characteristic curve (AUROC). Moreover, the performance of the proposed FSOS-AMC framework under low signal-to-noise ratio (SNR) conditions is much better than the compared schemes.
Index Terms:
Automatic modulation classification, few-shot open-set, multi-scale attention network, meta-prototype training, modular open-set classifier.
@@ I Introduction
Automatic modulation classification (AMC) is crucial for future wireless communication systems because it facilitates a range of applications including cognitive radio and intelligent spectrum management. AMC is the process of identifying the modulation scheme of a received signal, which is essential for both civilian and military applications. In civilian applications, AMC can be used for identifying unauthorized signals by detecting the modulation schemes, while in military applications, AMC facilitates abnormal signal detection and interference detection. With the further development of next-generation wireless communication systems, such as 6G, the AMC problem becomes more challenging due to the increasing number of modulation schemes and the dynamic wireless environment. Thus, it is essential to develop efficient and robust AMC frameworks.
Traditional AMC schemes include likelihood-based (LB) schemes and feature-based (FB) schemes [1, 2]. The LB schemes formulates AMC as a hypothesis testing problem, which requires known signal-to-noise ratio (SNR) and frequency offsets [3]. The FB methods are computationally expensive and require human-expert feature extraction [4]. Both methods are intricate and time-intensive, and they lack the flexibility required to adapt to the dynamic nature of wireless environments. In response to the rapid advancements in artificial intelligence [5, 6, 7], deep learning (DL)-based AMC frameworks have been proposed, which are capable of autonomously learning features directly from raw signals [8]. For example, the authors in [9] first developed a RadioML dataset and proposed a convolutional neural network (CNN)-based AMC framework for solving the AMC problem, achieving a high classification accuracy. Later, a large-scale dataset was proposed in [10] and a ResNet-based AMC framework was proposed for solving the AMC problem. However, the existing deep learning-based AMC methods relied on two dataset assumptions, large-scale training datasets and the same class pool between the training and testing data, i.e. a closed-set scenario. These assumptions failed under two scenarios, 1) few-shot conditions, where it is hard to collect enough labeled data and 2) open-set scenarios, where the testing classes do not appear during training.
Closed-set AMC frameworks: Closed-set is the most common scenario in AMC, where the training and testing data are from the same set of modulation schemes. Various deep learning-based AMC frameworks have been proposed, such as CNN-based [1], recurrent neural network (RNN)-based [11], and hybrid frameworks. The works [9] and [10] introduced the RadioML datasets, which are widely used for AMC research. Zhang et al. [1] proposed a novel multi-scale convolutional neural network-based AMC framework, which can achieve high classification accuracy. The authors in [11] exploited an RNN-based AMC framework, demonstrating that the RNN-based classifier is robust to uncertain noise conditions. In summary, while CNNs and RNNs have driven initial wireless deep learning research, a key limitation that remains is their reliance on large labeled datasets for training.
Few-shot AMC frameworks: few-shot AMC is a scenario where the model is trained with a few samples of each modulation scheme. Recent works can be categorized into three categories, namely, few-sample-based methods, support-data-based methods, and synthetic-data-based methods. Li et al. [12] introduced an AMR-CapsNet for AMC that achieves high accuracy under limited training data. The authors in [2] proposed a novel few-shot AMC framework by using semi-supervised learning, which can achieve high classification accuracy with a small number of samples. Existing few-shot AMC frameworks focused on the few-shot learning problem under the closed-set scenario, without considering the open-set recognition problem.
Open-set AMC frameworks: Open-set recognition (OSR) is a scenario where the model is trained with a set of modulation schemes, but the testing data may contain modulation schemes that are not seen during training. Bendale et al. [13] proposed the OpenMax model, replacing the softmax layer in the deep network with the OpenMax layer for OSR. Chen et al. [14] proposed a metric-based OSR framework, which can achieve high classification accuracy. The authors in [15] exploited a feature space singularity-based framework, which can improve discrimination between in-distribution and out-of-distribution features by promoting compact classification boundaries and reducing feature overlap. However, the existing open-set AMC frameworks still rely on large-scale training data.
Motivated by the challenge of the few-shot open-set AMC, we propose a novel few-shot open-set automatic modulation classification (FSOS-AMC) framework. The FSOS-AMC framework consists of three main parts, a multi-scale attention network, the meta-prototype training, and a modular open-set classifier. The multi-scale attention network can learn the multi-scale features of the input signal, the meta-prototype training is used to train the feature extractor in a few-shot learning manner, and the modular open-set classifier can classify the testing data into one of the known or potential unknown modulations. Simulation results demonstrate that the proposed FSOS-AMC framework can achieve high classification accuracy for known and unknown modulations compared to state-of-the-art methods. Moreover, the proposed FSOS-AMC framework can achieve high classification accuracy under low SNR conditions with less confusion between known and unknown modulations.
The structure of this paper is organized as follows. Section II introduces the preliminaries. Section III details our proposed FSOS-AMC framework. Section IV discusses the simulation results. The paper concludes in Section V.
@@ II-A Signal Model
Assume that the wireless communication system is equipped with a single transmitter, a channel and a receiver. The transmitter modulates the information bits into a signal, which is transmitted through the channel and received by the receiver. The received signal $x(n)$ can be denoted as
@@ III Proposed FSOS-AMC Framework
III-A Multi-Scale Attention Feature Extractor
Multi-scale module has been demonstrated useful for signal recognition tasks [1, 16, 17]. In this work, we introduce a multi-scale attention network (MSANet) to further enhance the feature representation. The MSANet consists of two layers of multi-scale attention modules, which are composed of a multi-scale module and a channel attention module. The multi-scale module is designed to capture the multi-scale features of the input signal, learning more separable features. Moreover, the channel attention mechanism is introduced to model the channel-wise dependencies and enhance the important features. Finally, a global average pooling layer is used to aggregate the feature maps and a $L_{2}$ normalization layer is used to normalize the feature vectors.
In the multi-scale module, a convolutional layer with kernel $3\times 1$, stride $2$ is used to downsample the input signal, labeling as $C_{\text{down}}$. Then, the downsampled signal is fed into four parallel convolutional layers with different kernel sizes. The four paralle convolutional layers consist of $1\times 1$, $3\times 1$, $5\times 1$, and $7\times 1$ kernels, respectively, and the convolutional operation is represented as $Conv_{1},Conv_{3},Conv_{5},Conv_{7}$. At the beginning of the $Conv_{3},Conv_{5},Conv_{7}$ layers, a $1\times 1$ convolutional layer is used to reduce the channel dimension, denoting as $C_{1}$. All the convolutional layers are followed by a batch normalization layer and a ReLU activation function. The output feature maps of the four convolutional layers are concatenated along the channel dimension. The multi-scale module can be formulated as
@@ III-A Multi-Scale Attention Feature Extractor
The excitation operation utilizes the channel descriptors generated in the squeeze operation to learn a channel-specific weight vector. This is typically implemented using a simple fully connected layer, including a ReLU activation followed by a Sigmoid activation to generate the weights. Mathematically, the excitation function can be represented as
@@
III-B Meta-Protype Training
As shown in Fig. 2, the feature extractor is trained using a supervised, episode-based methodology. The feature encoder MSANet, denoted as $f(\cdot)$, is systematically trained utilizing a supervised, episode-based methodology. Each episode commences with the data loader introducing a batch of training data procured from the main dataset. This batch is distinctively composed, ensuring an equal representation of samples from $N$ distinct classes.

### v2

@@ Abstract
Automatic modulation classification (AMC) plays a vital role in advancing future wireless communication networks. Although deep learning (DL)-based AMC frameworks have demonstrated remarkable classification capabilities, they typically require large-scale training datasets and assume consistent class distributions between training and testing data-prerequisites that prove challenging in few-shot and open-set scenarios. To address these limitations, we propose a novel few-shot open-set automatic modulation classification (FSOS-AMC) framework that integrates a multi-sequence multi-scale attention network (MS-MSANet), meta-prototype training, and a modular open-set classifier. The MS-MSANet extracts features from multi-sequence input signals, while meta-prototype training optimizes both the feature extractor and the modular open-set classifier, which can effectively categorize testing data into known modulation types or identify potential unknown modulations. Extensive simulation results demonstrate that our FSOS-AMC framework achieves superior performance in few-shot open-set scenarios compared to state-of-the-art methods. Specifically, the framework exhibits higher classification accuracy for both known and unknown modulations, as validated by improved accuracy and area under the receiver operating characteristic curve (AUROC) metrics. Moreover, the proposed framework demonstrates remarkable robustness under challenging low signal-to-noise ratio (SNR) conditions, significantly outperforming existing approaches.
@@ I Introduction
AUTOMATIC modulation classification (AMC) is crucial for future wireless communication systems because it facilitates a range of applications including cognitive radio and intelligent spectrum management. AMC is the process of identifying the modulation scheme of a received signal, which is essential for both civilian and military applications. In civilian applications, AMC can be used for identifying unauthorized signals by detecting the modulation schemes, while in military applications, AMC facilitates abnormal signal detection and interference detection. With the further development of next-generation wireless communication systems, such as sixth-generation (6G), the AMC problem becomes more challenging due to the increasing number of modulation schemes and the dynamic wireless environment. Thus, it is essential to develop efficient and robust AMC frameworks.
AMC aims at classifying the received signals into one of the known modulation schemes in non-cooperative communication environments without prior knowledge of the transmission parameters. AMC has numerous applications across the military and civilian domains, including decoding of intercepted transmissions, adaptive demodulation, and dynamic spectrum access (DSA). With the deep integration of artificial intelligence (AI) and future wireless communication, AMC has attracted extensive attention in the research community [1].
I-A Related Works and Motivations
@@
Traditional AMC schemes include likelihood-based (LB) schemes and feature-based (FB) schemes [2, 3]. The LB schemes formulate AMC as a hypothesis testing problem, which requires pre-known signal-to-noise ratio (SNR) and frequency offsets [4, 5]. The average likelihood ratio test (ALRT), the general likelihood ratio test (GLRT), hybrid likelihood ratio test (HLRT) methods are three main kinds of LB approaches. Although LB approaches, in theory, guarantee optimal classification results from the Bayesian sense, they have high computational complexity and require prior knowledge of channel parameters, which is not applicable in practice. FB methods have been proven to provide suboptimal solutions more efficiently due to their lower computational complexity [6]. The type of modulation is identified by extracting relevant features from the received signals. High-quality features are crucial for achieving robust performance with minimal complexity. Previous research has explored various features for AMC, including wavelet transforms [7], cyclostationary features [8], and higher-order cumulants [9]. In the classification phase, conventional classifiers such as random forests [10], $k$-nearest neighbors (KNN) [11], Naive Bayes [12], support vector machine (SVM) [13, 14], and multi-layer perception (MLP) [15] are used. However, these methods are often time-intensive as they rely on the extraction of handcrafted features, which necessitates substantial domain knowledge and engineering. However, the FB methods rely on the handcrafted features, which may not be the best representation for AMC. Moreover, the FB methods may not be suitable for the complex and dynamic wireless communication environment.
In response to the rapid advancements in AI, deep learning (DL) has been demonstrated powerful in solving various applications including computer vision [16, 17], geoscience [18, 19], as well as wireless communication [20, 2, 21]. DL-based AMC frameworks have been exploited, which can be capable of autonomously learning features directly from raw signals [22]. For example, the authors in [23] first developed a RadioML dataset and proposed a convolutional neural network (CNN)-based AMC framework for solving the AMC problem, achieving a high classification accuracy. Later, a large-scale dataset was proposed in [24] and a ResNet-based [25] AMC framework was proposed for solving the AMC problem. However, the existing deep learning-based AMC methods relied on two dataset assumptions, large-scale training datasets and the same class pool between the training and testing data, i.e. a closed-set scenario. These assumptions failed under two scenarios, 1) few-shot conditions, where it is hard to collect enough labeled data and 2) open-set scenarios, where the testing classes do not appear during training. While existing deep learning approaches have shown promise in modulation classification, their performance degrades significantly under realistic channel conditions, especially multipath fading. This challenge is particularly acute in few-shot and open-set scenarios where the model must generalize from limited training samples while dealing with channel distortions. The widely-used RadioML 2016.10A dataset, while valuable, inadequately represents authentic channel conditions, particularly the dynamic nature of propagation channels and severe multipath effects. This limitation has led to a gap between theoretical performance and practical deployability of existing AMC solutions.
Closed-set AMC frameworks: As shown in Fig. 1 (a), the closed-set scenario is the most common scenario in AMC, where the training and testing data are from the same set of modulation schemes. Various deep learning-based AMC frameworks have been proposed, such as CNN-based [2], recurrent neural network (RNN)-based [26], and hybrid frameworks. The works [23] and [24] introduced the RadioML datasets, which are widely used for AMC research. Zhang et al. [2] proposed a novel multi-scale convolutional neural network-based AMC framework, which can achieve high classification accuracy. The authors in [26] exploited an RNN-based AMC framework, demonstrating that RNN-based classifier is robust to the uncertain noise conditions. To solve the AMC problem under few-sample conditions, the authors in [3] proposed a novel few-shot AMC framework by using semi-supervised learning, which can achieve high classification accuracy with a few samples. In summary, while CNNs and RNNs have driven initial wireless deep learning research, a key limitation that remains is their reliance on large labeled datasets for training.
Signal representations: FB methods focus on extracting various features from the received signals. However, when DL methods are used, most of the existing works only use the IQ signal representation for several reasons. First, the IQ signal representation is the most common representation for AMC. Second, the IQ signal representation is easy to extract and can be directly fed into the DL model. However, the IQ signal representation may not be the best representation for AMC. The amplitude and phase (AP) representation and power spectral density (PSD) representation can provide more information about the signal, which may be useful for AMC. Qiu et al. [27] proposed DeepSIG, which is a hybrid heterogeneous modulation classification architecture. It combines RNN, CNN, and GNN models to process radio signals with multiple input formats, including IQ sequences, images mapped from IQ signals and graphs converted from IQ signals. The DeepSIG model can outperform single-input methods in classification accuracy, especially in few-shot scenarios. However, the combination of signal representations from images and graphs may not be the best representation for AMC for its complexity. Thus, it is essential to explore the best signal representation for AMC.
Channel effects: Deep learning models for modulation classification must be robust to diverse channel conditions in real-world applications. While the widely-used RadioML2016.10a dataset [23] has contributed significantly to the field, it inadequately represents authentic channel conditions, particularly the dynamic nature of propagation channels and severe multipath effects. Although the RadioML2018.01a dataset [24] attempted to address this limitation by incorporating over-the-air recordings of 24 digital and analog modulation types, its reliance on uncontrollable propagation environments makes it impossible to analyze the relationship between specific channel conditions and model performance. Additionally, being generated in a controlled laboratory environment, the dataset lacks diversity in crucial channel parameters such as fading and the number of taps. Building upon the foundation of RML2016.10a, the authors in [28] introduced RML22, which improves upon the original dataset’s methodology and data quality through a data-centric approach. Additionally, being generated in a controlled laboratory environment, the dataset lacks diversity in crucial channel parameters such as fading and the number of taps. To address these channel-related challenges, Wu et al. [29] developed a cumulant-based feature extraction method for AMC that achieves high classification accuracy under multipath fading channels. However, there remains a pressing need for datasets and models that can effectively handle diverse and realistic channel conditions.
Few-shot AMC frameworks: Few-shot AMC is a scenario where the model is trained with a few samples of each modulation scheme. Recent works can be categorized into three categories, namely, few-sample-based methods, support-data-based methods, and synthetic-data-based methods [3]. For few-sample-based methods, they only utilize the labeled data with few samples by using specifically designed architectures and transfer learning. Li et al. [30] introduced an AMR-CapsNet for AMC that achieves high accuracy under limited training data. The authors in [28] proposed a transfer learning method for AMC. The authors in [3] proposed a novel few-shot AMC framework by using semi-supervised learning, which can achieve high classification accuracy with a few samples. Existing few-shot AMC frameworks focused on the few-shot learning problem under the closed-set scenario, without considering the open-set recognition problem.
Open-set AMC frameworks: Open-set recognition (OSR) is a scenario where the model is trained with a set of modulation schemes, but the testing data may contain modulation schemes that are not seen during training. Bendale et al. [31] proposed the OpenMax model, replacing the softmax layer in the deep network with the OpenMax layer for OSR. Chen et al. [32] proposed a metric-based OSR framework, that can achieve high classification accuracy. The authors in [33] exploited a feature space singularity-based framework, which can improve discrimination between in-distribution and out-of-distribution features by promoting compact classification boundaries and reducing feature overlap. However, the existing open-set AMC frameworks still rely on large-scale training data.
As illustrated in Fig. 1 (d), few-shot open-set recognition represents a significantly more challenging paradigm compared to traditional closed-set, few-shot, and open-set recognition scenarios. This challenge becomes particularly relevant in automatic modulation classification (AMC) within complex and dynamic wireless communication environments, where systems must accurately classify signals using limited training samples while simultaneously identifying potential unknown modulation types. To address these practical challenges, we propose a novel few-shot open-set AMC framework that demonstrates robust classification performance for both known and unknown modulation types, even with limited training data.
I-B Contributions and Organization
Motivated by the challenges of few-shot open-set AMC under practical multipath fading channels, we propose a novel few-shot open-set automatic modulation classification (FSOS-AMC) framework. The FSOS-AMC framework consists of three main parts designed to address both limited training samples and channel impairments: a channel-aware multi-scale attention network, the meta-prototype training, and a modular open-set classifier. The multi-scale attention network adaptively learns robust multi-scale features from signals experiencing multipath fading, the meta-prototype training is used to train the feature extractor to be channel-invariant in a few-shot learning manner, and the modular open-set classifier enables accurate classification of testing data into known or potential unknown modulations even under severe channel conditions. Simulation results demonstrate that the proposed FSOS-AMC framework can achieve high classification accuracy for both known and unknown modulations compared to state-of-the-art methods across various multipath fading scenarios. Moreover, the proposed FSOS-AMC framework maintains robust performance under both low SNR conditions and challenging multipath environments with minimal confusion between known and unknown modulations.
  1. 1.
    A novel few-shot open-set automatic modulation classification (FSOS-AMC) framework is proposed to achieve robust classification for both known and unknown modulations under practical multipath fading channel conditions. The FSOS-AMC framework incorporates three key components: a channel-aware multi-scale attention network, meta-prototype training for few-shot learning, and a modular open-set classifier.
  2. 2.
    A multi-sequence multi-scale attention network (MS-MSANet) is developed to extract channel-robust features from multi-sequence input signals affected by multipath fading. The meta-prototype training is designed to learn channel-invariant features with limited training samples, while the modular open-set classifier enables accurate classification of both known and potential unknown modulations even under severe channel impairments.
  3. 3.
    Extensive simulation results under various multipath fading scenarios demonstrate the superiority of our proposed FSOS-AMC framework compared with other DL-based schemes in terms of classification accuracy for both known and unknown modulations. The framework maintains high performance even under challenging channel conditions. Moreover, the visualized features validate the framework’s ability to learn discriminative and channel-robust representations, highlighting its effectiveness in practical wireless environments.
The structure of this paper is organized as follows. Section II introduces the preliminaries, including the signal model and the problem formulation. Section III details our proposed FSOS-AMC framework, including the multi-sequence representation, the multi-sequence multi-scale attention network, meta-prototype training, and modular open-set classifier. Section IV discusses the simulation results. The paper concludes in Section V.
@@ II Preliminaries
In this section, the signal model is briefly introduced, and the signal preprocessing is implemented to acquire a good signal representation. Then, the problem formulation is presented.
@@ II-A Signal Model
Assume that the wireless communication system is equipped with a transmitter, a channel and a receiver. The transmitter modulates the information bits into a signal, which is transmitted through the channel and received by the receiver. The received signal $x(n)$ can be denoted as
@@ III Proposed FSOS-AMC Framework
In this section, we present the proposed few-shot open-set automatic modulation classification (FSOS-AMC) framework. Firstly, an overview of the proposed FSOS-AMC framework is presented. Then, the multi-sequence representation is introduced to acquire a good signal representation, followed by the multi-sequence multi-scale attention network (MS-MSANet) for feature extraction. Additionally, the meta-prototype training is adopted to train the feature extractor in a few-shot learning manner. Finally, the modular open-set classifier is used to classify the testing data into one of the known or potential unknown modulations.
III-A Overview of the Proposed Few-Shot Open-Set AMC (FSOS-AMC) Framework
@@
III-B Multi-Sequence Representations
The input signal can be represented in different ways, such as IQ signal representation, amplitude and phase (AP) representation, and power spectral density (PSD) representation, as shown in Fig. 3 and Fig. 4.
IQ signal representation
The IQ signal can be represented as in Eq. 2.
Amplitude and phase (AP) representation
The IQ signal representation can be converted into an amplitude ($A$) and phase ($P$) representation using the following formulas. Amplitude (A): This is calculated as the magnitude of the vector represented by the I and Q components. Mathematically, it is given by
$A=\sqrt{I^{2}+Q^{2}}.$ (4)
Phase ($P$) is the angle of the vector concerning the in-phase axis. It can be calculated using the arctangent function, specifically using the two-argument variant to take into account the correct quadrant of the angle. The formula is
$P=\arctan\left(\frac{Q}{I}\right).$ (5)
Power spectral density (PSD) representation
PSD representation is a fundamental tool in signal processing and communications, providing a detailed view of a signal’s power distribution across its frequency components. The transformation of IQ to PSD provides valuable insights into the frequency-domain characteristics of complex baseband signals. The PSD, $S_{xx}(f)$, of a wide-sense stationary random process $x(t)$ is formally defined as the Fourier transform of its autocorrelation function $R_{xx}(\tau)$
$S_{xx}(f)=\int_{-\infty}^{\infty}R_{xx}(\tau)e^{-j2\pi f\tau}d\tau.$ (6)
In practice, for discrete-time IQ signals, the PSD is often estimated using Welch’s method, which offers a balance between resolution and statistical variance. This method segments the signal, computes modified periodograms of these segments, and averages them to produce the PSD estimate
$\hat{P}{xx}(f)=\frac{1}{K}\sum{k=1}^{K}|\hat{X}_{k}(f)|^{2},$ (7)
where $K$ is the number of segments and $\hat{X}_{k}(f)$ is the discrete Fourier transform of the $k$-th windowed segment.
Figure 3: Different representations of the received signal of RadioML 2016.10A [23], including (a) IQ signal, (b) amplitude and phase (AP), and (c) power spectral density (PSD).
Figure 4: Different representations of the received signal of HisarMod2019.1 [34], including (a) IQ signal, (b) amplitude and phase (AP), and (c) power spectral density (PSD).
III-C Multi-Sequence Multi-Scale Attention Feature Extractor
The multi-sequence multi-scale attention feature extractor (MS-MSANet) is used to extract the features from the input signal. Multi-scale module has been demonstrated useful for signal recognition tasks [2, 35, 36]. In this work, we introduce a multi-scale attention network (MSANet) to further enhance the feature representation. The MSANet consists of two layers of multi-scale attention modules, which is composed of a multi-scale module and a channel attention module. The multi-scale module is designed to capture the multi-scale features of the input signal, learning more separable features. Moreover, the channel attention mechanism is introduced to model the channel-wise dependencies and enhance the important features. Finally, a global average pooling layer is used to aggregate the feature maps and an $L_{2}$ normalization layer is used to normalize the feature vectors.
In the multi-scale module, a convolutional layer with kernel $3\times 1$, stride $2$ is used to downsample the input signal, labeling it as $C_{\text{down}}$. Then, the downsampled signal is fed into four parallel convolutional layers with different kernel sizes. The four parallel convolutional layers consist of $1\times 1$, $3\times 1$, $5\times 1$, and $7\times 1$ kernels, respectively, and the convolutional operation is represented as $Conv_{1},Conv_{3},Conv_{5},Conv_{7}$. At the beginning of the $Conv_{3},Conv_{5},Conv_{7}$ layers, a $1\times 1$ convolutional layer is used to reduce the channel dimension, denoting $C_{1}$. All the convolutional layers are followed by a batch normalization layer and a ReLU activation function. The output feature maps of the four convolutional layers are concatenated along the channel dimension. The multi-scale module can be formulated as
@@ III-C Multi-Sequence Multi-Scale Attention Feature Extractor
The excitation operation utilizes the channel descriptors generated in the squeeze operation to learn a channel-specific weight vector. This is typically implemented using a simple fully connected layer, including an ReLU activation followed by a Sigmoid activation to generate the weights. Mathematically, the excitation function can be represented as
@@
III-D Meta-Protype Training
As shown in Fig. 2, the feature extractor is trained using a supervised, episode-based methodology. The feature encoder MSANet, denoted as $f(\cdot)$, is systematically trained utilizing a supervised, episode-based methodology. Each episode commences with the data loader introducing a batch of training data procured from the main dataset. This batch is distinctively composed, ensuring an equitable representation of samples from $N$ distinct classes.